What is the equation of line which is pararrel to (3x-2y=5)and passes through the point (-2,-4)
1
Expert's answer
2019-10-28T12:50:08-0400
  1. We transform the general equation of a given line into a slope-intercept equation: 3x-2y=5; -2y=-3x+5; y=1.5x-2.5
  2. The condition of parallel straight lines is k1=k2. Then we obtain y=1.5x+b.
  3. A new line goes through the point (-2; -4):

-4=1.5*(-2)+b; -4=-3+b; b=-4+3; b=-1.

Consequently, the equation of the required straight line is y=1.5x-1, or 2y=3x-2, or 3x-2y=2.